# Thermoelectric processes of quantum normal-superconductor interfaces

### v1

@@ III.1 Breaking energy symmetry in Coulomb blockaded quantum dots
In this section we discuss, even in the presence o superconducting proximity effects, how to break EIS. To do that it is important to include in the description the tunneling through the contacts to normal reservoirs $l$ which corresponds to a transition process between even and odd parity (number of electrons) states. The tunneling term is
@@
where $\hat{\delta}_{\sigma}^{e}=\hat{d}_{\sigma}^{\dagger}$, $\hat{\delta}_{\sigma}^{h}=\hat{d}_{\sigma}$ and $f_{l}^{\alpha}(E)$ is the Fermi function for the normal reservoir with density of states $\nu_{l}$ at electrochemical potential $\mu_{l}$ and temperature $T_{l}$. Clearly, $|\langle\sigma|\hat{d}_{\sigma}^{\dagger}|\pm\rangle|^{2}\propto{A}_{\pm}^{2}$ is different from $|\langle\sigma|\hat{d}_{\sigma}|\pm\rangle|^{2}\propto\Gamma_{S}^{2}$, except for the electron-hole symmetry point $\tilde{\varepsilon}=0$. Hence, the electron/hole contribution of each transition can be controlled by tuning the quantum dot gate voltage. This will also affect the occupation of the different states.
@@
The regime dominated by quasiparticles at higher temperatures or weak coupling to the superconductor offers a clearer picture. There, ignoring the Andreev reflection contributions, the particle tunneling to the superconductor is governed by the BCS density of states of Eq. (11). The tunneling rates will be proportional to $\nu_{S}^{n}=\nu_{S}(\varepsilon+nU)$, where $n$ is the minimum occupation of the quantum dot before or after the process, see Fig. 2(b). The effect of the gap in $\nu_{S}^{n}$ prevents charge to flow for quantum dot energies around the chemical potential except when $\Delta\leq U/2$. In a sequential tunneling description, it reads $I\propto\sum_{nn^{\prime}}(f_{Nn}+\nu_{S}^{n}f_{Sn})\nu_{S}^{n^{\prime}}(f_{Nn^{\prime}}-f_{Sn^{\prime}})(1-\delta_{nn^{\prime}})$, where $f_{ln}=f_{l}(\varepsilon+nU)$, see Fig. 2(c). The current is hole(electron)-like for $\tilde{\varepsilon}<0(>0)$, as would correspond to an all-normal configuration [138], with sharp features corresponding to the peaked density of states at the gap borders.
@@
When the superconducting terminals are more than one it is even possible to establish also a phase differences between superconducting terminals. In such case it is possible to investigate phase-dependent effects on heat and charge transport and thermoelectrics [155, 156, 157, 158]. Intriguing thermoelectrical effects are also possible when superconductors and quantum dots are coupled with ferromagnetic materials [159, 142, 160]. Nonlocal cooling mechanism in ballistic Andreev interferometers have also been proposed recently [161].
@@ III.3 Thermoelectricity by Cooper pair splitting
This geometry can be used to define a nonlocal thermoelectric engine: a temperature difference between the two normal reservoirs generates a current out of the superconductor [184, 35, 36, 185, 186]. Experiments have confirmed this mechanism [187]. When the quantum dot level energies are antisymmetric around the superconductor electrochemical potential, energy-conserving Cooper pair splitting processes are privileged over EC, such that the same current flows ongoing (outgoing) from (to) the two normal leads into (from) the superconducting terminal [35, 36]. Heat is hence converted into positively cross-correlated charge flows [188] and can also lead to coherent Peltier cooling, which serves as a prove of Cooper pair splitting [35]. Differently EC contributes to the local thermoelectric response between the normal reservoirs [35, 36, 185], in a similar way to what one expects from all-semiconductor double quantum dots [189, 190]. Configurations with EC and CAR having similar contributions can be beneficial for enhancing power and efficiency when the hot terminal is a voltage probe [35] or determining strong nonlocal responses from heating effects [191].
@@ III.4 Detection of Majoranas
Since the early investigations into Majorana states in condensed matter systems, thermoelectric effects have been identified as a potential signature of their presence. Majorana signatures have been pointed out in thermal conductance [192, 157], voltage thermopower [193, 194], or the violation of the Wiedemann-Franz law [195, 44]. In quantum-dot-based multiterminal setups, it has been suggested that a key signature for detecting Majorana zero modes (MZMs) is the opposite sign behavior of the Seebeck coefficient as a function of the quantum dot energy level, in contrast to what is observed in conventional superconductors [42, 43, 196, 197, 198, 199, 200, 201]. Most of these works just consider the effect of the MZMs and neglect the contribution of quasiparticle states above the gap, whose effect in the out of equilibrium response of these systems has remained largely unexplored [202, 203] but can be included by means of Green function techniques [204]. The thermoelectric response of more complex geometries including a double QD connected in series [205] or in parallel [45] with a topological superconductor hosting MZMs have also been explored.
@@ IV.1 Helical edge states
In Ref. 218, it was shown that a two-terminal device of a 2DTI hosting a magnetic island is an optimal thermoelectric heat engine. Here, we briefly review the thermoelectric properties of the device studied in Refs.  [40, 41], where the 2DTI is contacted to superconductors forming a Josephson junction, cf. Fig. 3(a). The edge modes are contacted by left ($S_{L}$) and right ($S_{R}$) $s$-wave type superconductors. Considering the Nambu basis $\psi^{T}=\left(\psi_{\uparrow},\psi_{\downarrow},\psi^{\dagger}_{\downarrow},-\psi^{\dagger}_{\uparrow}\right)$, the BdG Hamiltonian now becomes
@@
where $\Delta(x)=\Delta\left[\theta(-x-L/2)e^{i\phi_{L}}+\theta(x-L/2)e^{i\phi_{R}}\right]$ is the pairing potential induced on the edge modes within the region where they are proximized to the superconducting contacts. A phase bias $\phi=\phi_{L}-\phi_{R}$ is considered between the two superconductors, which define real and imaginary parts of the pairing potential ($\Delta^{\prime}(x)$ and $\Delta^{\prime\prime}(x)$, respectively). The other ingredient is a Doppler shift [219] $p_{DS}=(\pi\hbar/L)\Phi/\Phi_{0}$ generated by the magnetic flux $\Phi$ within the Josephson junction region of TI between the superconductors, i.e., $-L/2\leq x\leq L/2$, with $\Phi_{0}$ the magnetic flux quantum. This Hamiltonian is particle-hole symmetric, and the naive expectation for the symmetric right-left configuration is the absence of any thermoelectric local response between the superconducting terminals when a temperature bias is imposed at the superconductors. However, in Refs. [40, 41], it was shown that nonlocal thermoelectricity is still possible in such a system by adding a third normal-metal terminal, see Fig. 3(a).
@@
The mechanism generating the nonlocal thermoelectric response is understood by analyzing the spectrum of the full Hamiltonian, shown in Fig. 3(b). We can identify the particle-hole symmetry of the BdG Hamiltonian, with positive and negative energies related by $\varepsilon_{k}=-\varepsilon_{-k}$. In the different branches of the spectrum, the electron-like (e) and hole-like (h) character of the quasiparticles is indicated with solid and dashed lines, respectively. We can identify two cones. Each one has left (--) and right ($+$) moving quasiparticles (plotted in different colors). The Doppler shift breaks the symmetry under the inversion $k\leftrightarrow-k$ of the spectrum. As a consequence, the main contribution to the transport window is originated in a single cone hosting states of quasielectrons and quasiholes traveling in opposite directions. Importantly for our purposes, however, the two types of particles are injected in different reservoirs at different temperatures. Since they come from opposite thermal sources they introduce an extra sign when the current is computed in the normal middle terminal such as ${\cal I}_{N}\propto\sum_{\alpha}\alpha({\cal T}^{e\alpha}_{NL}-{\cal T}^{e\alpha}_{NR})$ with respect to the local term ${\cal I}_{R}\propto\sum_{\alpha}\alpha({\cal T}^{e\alpha}_{RL}+{\cal T}^{e\alpha}_{LR})$. Therefore, even if the local thermoelectric response is null a net particle charge flux is generated at the normal probe (nonlocal response) [39], as sketched in Fig. 3(c). The reciprocal process injecting a net charge current at the normal probe is sketched in Fig. 3(d). Such a flux is converted into uncompensated fluxes of quasielectrons and quasiholes in the proximized TI, generating a net heat flux between the two superconducting contacts.
@@
[…]
@@
Another related proposal is based on a thermally biased Andreev interferometer with a middle ferromagnetic island on top of the helical edge of a 2DTI [224]. Without the magnetic impurity, the thermoelectric conversion is analyzed in Ref. 225. In the presence of the magnetic domain, a spin-polarized thermal supercurrent emerges between the superconductors due to a unique interference effect of CARs, that is absent for EC processes. The resulting thermoelectric effect can be turned on and off by tuning the phase difference between the superconducting leads and does not rely on manipulating the ferromagnetic domain.
@@ IV.2 Superconducting helical wires
A similar effect to the one discussed in the previous section is predicted to take place in semiconducting wires with Rashba spin orbit coupling (SOC) proximized to superconductors and under the effect of a magnetic field [39]. These types of structures are used to implement the 1D topological superconducting phases hosting Majorana edge modes experimentally [226, 227], motivated by the model proposed in Refs. 228, 42. Experimentally, these wires are fabricated with GaAs doped with In and proximized with a $s$-wave superconductor of gap $\Delta$ [229, 230, 231, 232, 233, 234]. The Hamiltonian reads
@@
where $\lambda_{k}=\alpha k$ represents the SOC acting on the spin degrees of freedom along the direction $\vec{n}_{\lambda}$ and $\Delta_{B}=g\mu_{B}B$ is the Zeeman term associated to the magnetic field $B$ acting along the direction $\vec{n}_{B}$. $\xi_{k}=k^{2}/2m-\mu$ is the kinetic energy relative to the chemical potential $\mu$. When such wires are contacted at the ends with reservoirs at different temperatures $T\pm\delta T/2$ and contacted at an intermediate point with a normal conducting probe at temperate $T$, see Fig. 4(a), a nonlocal thermoelectric response is predicted when the angle between the magnetic field and the SOC departs from perpendicular orientations, $\theta=\cos^{-1}(\vec{n}_{B}\cdot\vec{n}_{\lambda})\neq n\pi/2$.
@@
The origin of the nonlocal response can be understood by analyzing the spectrum. Without superconductivity, the spectrum contains two bands and a gap defined by the combined effect of the SOC and the Zeeman field. Adding superconductivity and expressing the Hamiltonian in Nambu basis leads to the spectra shown in Fig. 4(b). The relevant situation for thermoelectric response takes place when the chemical potential is close to the bottom of the lower band, precisely the case illustrated in the first and last panels of Fig. 4(b). It is important to notice that when the magnetic field and the SOC are exactly perpendicular [see central panel in Fig. 4(b)] the spectrum is fully symmetric with respect to changing $k\rightarrow-k$. Instead, for other angles, one of the quasiparticle cones with positive energy moves to lower energies and eventually crosses zero, defining two Bogoliubov Fermi points. This situation is identified with a gray dashed-line box in Fig. 4(b) and is similar to the one discussed in the analysis of Fig. 3. In fact, the low-energy sector of the spectrum is composed of a single branch of quasielectrons moving to the right and quasiholes moving to the left. These uncompensated fluxes generate an electron current at the normal lead described in linear response by Eq. (20). Instead, for higher values of the chemical potential, the two bands of the model without superconductivity contribute to the transport processes. In this situation, there are a pair of left- and right-moving quasielectron as well as a pair of left- and right-moving quasihole branches. The net charge flux is compensated in this case and the electron current in the normal probe is vanishing small. Interestingly, the range of chemical potentials for the topological phase to take place is $\mu^{2}\leq\Delta_{B}^{2}-\Delta^{2}$ and coincides with the one for which a non-vanishing thermoelectric response is expected. It is also interesting that the range of angles for which the topological phase takes place is complementary to the one with important thermoelectric response. Namely, the topological phase is predicted within $|\theta-\pi/2|\leq\delta_{c}$, while the thermoelectric response is predicted within $|\theta-\pi/2|>\delta_{c}$. Note that this nonlocal thermoelectric effect is not a signature of Majorana fermions per sé. Nevertheless, it provides very relevant information to characterize the topological phase and the relative alignment between the SOC and the magnetic field to identify where in the wide parameter space the topological transition should localize. The nonlocal thermoelectric transport requires a proximized helical system in the electron bands which are a necessary condition for the nontrivial topological phase.

### v2

@@ III.1 Breaking energy symmetry in Coulomb blockaded quantum dots
In this section we discuss, even in the presence of superconducting proximity effects, how to break EIS. To do that it is important to include in the description the tunneling through the contacts to normal reservoirs $l$ which corresponds to a transition process between even and odd parity (number of electrons) states. The tunneling term is
@@
where $\hat{\delta}_{\sigma}^{e}=\hat{d}_{\sigma}^{\dagger}$, $\hat{\delta}_{\sigma}^{h}=\hat{d}_{\sigma}$ and $f_{l}^{\alpha}(E)$ is the Fermi function for the normal reservoir with density of states $\nu_{l}$ at electrochemical potential $\mu_{l}$ and temperature $T_{l}$. Clearly, $|\langle\sigma|\hat{d}_{\sigma}^{\dagger}|\pm\rangle|^{2}\propto{A}_{\pm}^{2}$ is different from $|\langle\sigma|\hat{d}_{\sigma}|\pm\rangle|^{2}\propto\Gamma_{S}^{2}$, except for the electron-hole symmetry point $\tilde{\varepsilon}=0$. Hence, the electron/hole contribution of each transition can be controlled by tuning the quantum dot gate voltage. This will also affect the occupation of the different states. In the sequential tunneling regime, one uses these rates to write a master equation $\dot{\rho}_{\kappa\kappa}=\sum_{l\alpha\lambda}(\gamma_{\kappa\lambda}^{l\alpha}\rho_{\lambda\lambda}-\gamma_{\lambda\kappa}^{l\alpha}\rho_{\kappa\kappa})$ for the state occupations, $\rho_{\kappa\kappa}$ [138], and writes the particle and heat currents out of terminal $l$ [142]
$I_{l}$ $=\underset{\lambda,\kappa}{\sum}\left(\gamma_{\lambda\kappa}^{le}-\gamma_{\lambda\kappa}^{lh}\right)\rho_{\kappa\kappa},$ (20)
$J_{l}$ $=\underset{\lambda,\kappa}{\sum}\left(E_{\lambda}-E_{\kappa}-\mu_{l}\right)\left(\gamma_{\kappa\lambda}^{le}\rho_{\lambda\lambda}-\gamma_{\lambda\kappa}^{lh}\rho_{\kappa\kappa}\right).$ (21)
using the steady state solution, $\dot{\rho}=0$.
@@
The regime dominated by quasiparticles at higher temperatures or weak coupling to the superconductor offers a clearer picture. There, ignoring the Andreev reflection contributions, the particle tunneling to the superconductor is governed by the BCS density of states of Eq. (13). The tunneling rates will be proportional to $\nu_{S}^{n}=\nu_{S}(\varepsilon+nU)$, where $n$ is the minimum occupation of the quantum dot before or after the process, see Fig. 2(b). The effect of the gap in $\nu_{S}^{n}$ prevents charge from flowing for quantum dot energies around the chemical potential except when $\Delta\leq U/2$. In a sequential tunneling description, it reads $I\propto\sum_{nn^{\prime}}(f_{Nn}+\nu_{S}^{n}f_{Sn})\nu_{S}^{n^{\prime}}(f_{Nn^{\prime}}-f_{Sn^{\prime}})(1-\delta_{nn^{\prime}})$, where $f_{ln}=f_{l}(\varepsilon+nU)$, see Fig. 2(c). The current is hole(electron)-like for $\tilde{\varepsilon}<0(>0)$, as would correspond to an all-normal configuration [144], with sharp features corresponding to the peaked density of states at the gap borders.
@@
When the superconducting terminals are more than one it is als possible to establish phase differences between superconducting terminals. In such case it is possible to investigate phase-dependent effects on heat and charge transport and thermoelectrics [161, 162, 163, 164]. Intriguing thermoelectrical effects are also possible when superconductors and quantum dots are coupled with ferromagnetic materials [165, 148, 166]. Nonlocal cooling mechanism in ballistic Andreev interferometers have also been proposed recently [167].
@@ III.3 Thermoelectricity by Cooper pair splitting
This geometry can be used to define a nonlocal thermoelectric engine: a temperature difference between the two normal reservoirs generates a current out of the superconductor [190, 35, 36, 191, 192]. Experiments have confirmed this mechanism [193]. When the quantum dot level energies are antisymmetric around the superconductor electrochemical potential, energy-conserving Cooper pair splitting processes are privileged over EC, such that the same current flows ingoing (outgoing) from (to) the two normal leads into (from) the superconducting terminal [35, 36]. This nonlocal splitting explicitly breaks the EIS such that, e.g., ${\cal T}_{HC}^{eh}(E)\neq{\cal T}_{HC}^{eh}(-E)$. Heat is hence converted into positively cross-correlated charge flows [194] and can also lead to coherent Peltier cooling, which serves as a prove of Cooper pair splitting [35]. Differently EC contributes to the local thermoelectric response between the normal reservoirs [35, 36, 191], in a similar way to what one expects from all-semiconductor double quantum dots [195, 196]. Configurations with EC and CAR having similar contributions can be beneficial for enhancing power and efficiency when the hot terminal is a voltage probe [35] or determining strong nonlocal responses from heating effects [197].
@@ III.4 Detection of Majoranas
Since the early investigations into Majorana states in condensed matter systems, thermoelectric effects have been identified as a potential signature of their presence. Majorana signatures have been pointed out in thermal conductance [198, 163], voltage thermopower [199, 200], or the violation of the Wiedemann-Franz law [201, 44]. In quantum-dot-based multiterminal setups, it has been suggested that a key signature for detecting Majorana zero modes (MZMs) is the opposite sign behavior of the Seebeck coefficient as a function of the quantum dot energy level, in contrast to what is observed in conventional superconductors [42, 43, 202, 203, 204, 205, 206, 207]. Most of these works just consider the effect of the MZMs and neglect the contribution of quasiparticle states above the gap, whose effect in the out of equilibrium response of these systems has remained largely unexplored [208, 209] but can be included by means of Green function techniques [210]. The thermoelectric response of more complex geometries including a double QD connected in series [211] or in parallel [45] with a topological superconductor hosting MZMs has also been explored.
@@ IV.1 Helical edge states
In Ref. 224, it was shown that a two-terminal device of a 2DTI hosting a magnetic island is an optimal thermoelectric heat engine. Here, we briefly review the thermoelectric properties of the device studied in Refs.  [40, 41], where the 2DTI is contacted to superconductors forming a Josephson junction, cf. Fig. 3(a). The edge modes are contacted by left ($S_{L}$) and right ($S_{R}$) $s$-wave type superconductors. Considering the Nambu basis $\psi^{T}=\left(\psi_{\uparrow},\psi_{\downarrow},\psi^{\dagger}_{\downarrow},-\psi^{\dagger}_{\uparrow}\right)$, the BdG Hamiltonian now becomes
@@
where $\Delta(x)=\Delta\left[\theta(-x-L/2)e^{i\phi_{L}}+\theta(x-L/2)e^{i\phi_{R}}\right]$ is the pairing potential induced on the edge modes within the region where they are proximized to the superconducting contacts. A phase bias $\phi=\phi_{L}-\phi_{R}$ is considered between the two superconductors, which define real and imaginary parts of the pairing potential ($\Delta^{\prime}(x)$ and $\Delta^{\prime\prime}(x)$, respectively). The other ingredient is a Doppler shift [225] $p_{DS}=(\pi\hbar/L)\Phi/\Phi_{0}$ generated by the magnetic flux $\Phi$ within the Josephson junction region of TI between the superconductors, i.e., $L/2\leq x\leq L/2$, with $\Phi_{0}$ the magnetic flux quantum. This Hamiltonian is particle-hole symmetric, and the naive expectation for the symmetric right-left configuration is the absence of any thermoelectric local response between the superconducting terminals when a temperature bias is imposed at the superconductors. However, in Refs. [40, 41], it was shown that nonlocal thermoelectricity is still possible in such a system by adding a third normal-metal terminal, see Fig. 3(a).
@@
The mechanism generating the nonlocal thermoelectric response is understood by analyzing the spectrum of the full Hamiltonian, shown in Fig. 3(b). We can identify the particle-hole symmetry of the BdG Hamiltonian, with positive and negative energies related by $\varepsilon_{k}=-\varepsilon_{-k}$. In the different branches of the spectrum, the electron-like (e) and hole-like (h) character of the quasiparticles is indicated with solid and dashed lines, respectively. We can identify two cones. Each one has left (-) and right (++) moving quasiparticles (plotted in different colors). The Doppler shift breaks the symmetry under the inversion $k\leftrightarrow-k$ of the spectrum. As a consequence, the main contribution to the transport window originates in a single cone hosting states of quasielectrons and quasiholes traveling in opposite directions. Importantly for our purposes, however, the two types of particles are injected in different reservoirs at different temperatures. Since they come from opposite thermal sources they introduce an extra sign when the current is computed in the normal middle terminal such as ${\cal I}_{N}\propto\sum_{\alpha}\alpha({\cal T}^{e\alpha}_{NL}-{\cal T}^{e\alpha}_{NR})$ with respect to the local term ${\cal I}_{R}\propto\sum_{\alpha}\alpha({\cal T}^{e\alpha}_{RL}+{\cal T}^{e\alpha}_{LR})$. Therefore, even if the local thermoelectric response is null a net particle charge flux is generated at the normal probe (nonlocal response) [39], as sketched in Fig. 3(c). The reciprocal process, injecting a net charge current at the normal probe, is sketched in Fig. 3(d). Such a flux is converted into uncompensated fluxes of quasielectrons and quasiholes in the proximized TI, generating a net heat flux between the two superconducting contacts.
@@
$\begin{aligned} J^{0}_{N}&=L^{(nl)}_{11}\left(V_{N}/T\right)+L_{12}^{(nl)}\left(\delta T/T^{2}\right),\\ J^{0}_{N}&=L_{21}^{(nl)}\left(V_{N}/T\right)+L_{22}^{(nl)}\left(\delta T/T^{2}\right),\end{aligned}$ (24)
@@
Another related proposal is based on a thermally biased Andreev interferometer with a middle ferromagnetic island on top of the helical edge of a 2DTI [230]. Without the magnetic impurity, the thermoelectric conversion is analyzed in Ref. 231. In the presence of the magnetic domain, a spin-polarized thermal supercurrent emerges between the superconductors due to a unique interference effect of CARs, that is absent for EC processes. The resulting thermoelectric effect can be turned on and off by tuning the phase difference between the superconducting leads and does not rely on manipulating the ferromagnetic domain.
@@ IV.2 Superconducting helical wires
A similar effect to the one discussed in the previous section is predicted to take place in semiconducting wires with Rashba spin-orbit coupling (SOC) proximized to superconductors and under the effect of a magnetic field [39]. These types of structures are used to implement the 1D topological superconducting phases hosting Majorana edge modes experimentally [232, 233], motivated by the model proposed in Refs. 234, 42. Experimentally, these wires are fabricated with GaAs doped with In and proximized with a $s$-wave superconductor of gap $\Delta$ [235, 236, 237, 238, 239, 240]. The Hamiltonian reads
@@
where $\lambda_{k}=\alpha k$ represents the SOC acting on the spin degrees of freedom along the direction $\vec{n}_{\lambda}$ and $\Delta_{B}=g\mu_{B}B$ is the Zeeman term associated to the magnetic field $B$ acting along the direction $\vec{n}_{B}$. $\xi_{k}=k^{2}/2m-\mu$ is the kinetic energy relative to the chemical potential $\mu$. When such wires are contacted at the ends with reservoirs at different temperatures $T\pm\delta T/2$ and contacted at an intermediate point with a normal conducting probe at temperature $T$, see Fig. 4(a), a nonlocal thermoelectric response is predicted when the angle between the magnetic field and the SOC departs from perpendicular orientations, $\theta=\cos^{-1}(\vec{n}_{B}\cdot\vec{n}_{\lambda})\neq n\pi/2$.
@@
The origin of the nonlocal response can be understood by analyzing the spectrum. Without superconductivity, the spectrum contains two bands and a gap defined by the combined effect of the SOC and the Zeeman field. Adding superconductivity and expressing the Hamiltonian in Nambu basis leads to the spectra shown in Fig. 4(b). The relevant situation for thermoelectric response takes place when the chemical potential is close to the bottom of the lower band, precisely the case illustrated in the first and last panels of Fig. 4(b). It is important to notice that when the magnetic field and the SOC are exactly perpendicular [see central panel in Fig. 4(b)] the spectrum is fully symmetric with respect to changing $k\rightarrow-k$. Instead, for other angles, one of the quasiparticle cones with positive energy moves to lower energies and eventually crosses zero, defining two Bogoliubov Fermi points. This situation is identified with a gray dashed-line box in Fig. 4(b) and is similar to the one discussed in the analysis of Fig. 3. In fact, the low-energy sector of the spectrum is composed of a single branch of quasielectrons moving to the right and quasiholes moving to the left. These uncompensated fluxes generate an electron current at the normal lead described in linear response by Eq. (24). Instead, for higher values of the chemical potential, the two bands of the model without superconductivity contribute to the transport processes. In this situation, there are a pair of left- and right-moving quasielectrons as well as a pair of left- and right-moving quasihole branches. The net charge flux is compensated in this case and the electron current in the normal probe is vanishingly small. Interestingly, the range of chemical potentials for the topological phase to take place is $\mu^{2}\leq\Delta_{B}^{2}-\Delta^{2}$ and coincides with the one for which a non-vanishing thermoelectric response is expected. It is also interesting that the range of angles for which the topological phase takes place is complementary to the one with important thermoelectric response. Namely, the topological phase is predicted within $|\theta-\pi/2|\leq\delta_{c}$, while the thermoelectric response is predicted within $|\theta-\pi/2|>\delta_{c}$. Note that this nonlocal thermoelectric effect is not a signature of Majorana fermions per se. Nevertheless, it provides very relevant information to characterize the topological phase and the relative alignment between the SOC and the magnetic field to identify where in the wide parameter space the topological transition should localize. The nonlocal thermoelectric transport requires a proximized helical system in the electron bands, which is also a necessary condition for the nontrivial topological phase.